A torque of 0.97 NM is applied to a bicycle wheel of radius 45 cm and mass 0.80 kg.
Treating the wheel as a hoop, find its angular acceleration.
Answer #1

τ = mr^2α

α = τ/mr^2 = 0.97/(0.8*0.45^2) = 5.988 rad/s^2